Steiner Corporation manufactures water skis through two processes: Molding and Packaging. In the Molding Department, fiberglass is heated and shaped into the form of a ski. In the Packaging Department, the skis are placed in cartons and sent to the finished goods warehouse. Materials are entered at the beginning of both processes. Labor and manufacturing overhead are incurred uniformly throughout each process. Production and cost data for the Molding Department for January 2017 are presented below.

Production Data

January

Beginning work in process units

–0–

Units started into production

50,000

Ending work in process units

2,500

Percent complete - ending inventory

40%

Cost Data

January

Materials

$510,000

Labor

92,500

Overhead

  150,000

   Total

$752,500

Instructions

(a)  Compute the physical units of production.

(b) Determine the equivalent units of production for materials and conversion costs.

(c)  Compute the unit costs of production.

(d) Determine the costs to be assigned to the units transferred out and in process.

(a)

Physical Units:

Units to be accounted for:

Work in process, January 1

Started into productiom

          Total Units

Units accounted for:

Transferred out

Work in process, January 31

          Total Units

(b)

Equivalent Units:

Materials

Conversion Costs

Units transferred out

Work in Process Jan 31

Total Equivalent Units

(c)

Unit Costs

Materials

Conversion cost

Total unit cost

(d)

Costs accounted for:

Transferred out

Work in process, Jan 31:

Materials

Conversion costs

          Total costs
Answer #1

(a)

Physical Units:

Units to be accounted for:

Work in process, January 1

0

Started into productiom

50000

          Total Units

50000

Units accounted for:

Transferred out

47500

Work in process, January 31

2500

          Total Units

50000

(b)

Equivalent Units:

Materials

Conversion Costs

Units transferred out

47500 47500

Work in Process Jan 31

Material 2500*100% = 2500
Conversion 2500*40% = 1000

Total Equivalent Units

50000 48500

(c)

Unit Costs

Materials

510000/50000 = 10.20

Conversion cost

242500/48500 = 5

Total unit cost

15.20

(d)

Costs accounted for:

Transferred out (47500*15.20)

722000

Work in process, Jan 31:

Materials (2500*10.20)

25500

Conversion costs (1000*5)

5000 30500

          Total costs

752500